Grade upto college level Physical Chemistry

The average velocity of gas molecules is 400 m/sec. Calculate its rms velocity at the same temperature.

Shane Macguire
12 Years agoGrade upto college level

2 Answers

Deepak Patra
12 Years ago
Crms = √3 RT/M , Cav √8 RT/πM
Crms/Cav = √3 RT/M * √πM/8 RT = √3π/8 = 1.085
Crms = 1.085 * Cav = 1.085 * 400 = 434 ms-1
Sparsh
8 Years ago
Use the formula of Crms and Cav and then divide both of them and then the will come approximately 434 meter per second